OpenStudy (anonymous):

For what value of k does the equation e^(2x)=k*sqrt(x) have exactly one solution?

OpenStudy (amistre64):

wed have to scale sqrt(x) to reach i spose

OpenStudy (amistre64):

somehow

OpenStudy (amistre64):

e^(2x)=k*sqrt(x) 2x = ln(k*sqrt(x)) 2x = ln(k) + ln(sqrt(x)) 2x-ln(sqrt(x)) = ln(k) maybe in the right direction?

OpenStudy (anonymous):

im nt so sure but that wht i have k has to be bigger than zero, because exp(2x)>0 en sqrt(x) > 0. Second : x >= 0, because sqrt(x) only defined for real values x >= 0 For x >= 0, exp(2x) >= 1, so k sqrt(x) >= 1 Take x = 0, then we have 1 = k * 0, so k has to be +infinity Take x = 0.5 then we have e = k / sqrt(2) => k = e sqrt(2) Take x = 1, then we have e² = k, so k has to be e² Take x = 2, then we have e^4 = k sqrt(2), so k has to be e^4 / sqrt(2) So we have a minimum for k. We have to calculate that minimum k = exp(2x) / sqrt(x) => k ' = ( 2 exp(2x) sqrt(x) - exp(2x) / 2 sqrt(x) ) / x = 0 2 sqrt(x) - 1/2 sqrt(x) = 0 => 4 x - 1 = 0 => x = 1/4 if x = 1/4, we have exp(0.5) = k / 2 => k = 2 exp(0.5) So we have exactly one solution for k = 2 exp(0.5) and zero solutions for k < 2 exp(0.5) and two solutions for k > 2 exp(0.5)

OpenStudy (amistre64):

e^(2x-ln(sqrt(x))) = k e^(2x) * e^(-ln(sqrt(x))) = k

OpenStudy (anonymous):

im nt so sure tho

OpenStudy (amistre64):

e^(2x)-k*sqrt(x)=0 i cant be sure either but, i know its between sqrt(10.8) and sqrt(10.9) :)

OpenStudy (amistre64):

i think the wolf says your right http://www.wolframalpha.com/input/?i=e^%282x%29-%282+exp%280.5%29%29*sqrt%28x%29%3D0